OpenStudy (brucebaner):

The sum of the measures of the interior angles of a polygon is 720°. How many sides does the polygon have?

OpenStudy (ajanijones):

when the sum of interior angles is 720, therefore, (n-2)*180 = 720 n-2 = 720/180 n-2 = 4 n = 6 Medal :O